paul is standing at the top of an escalator and lily is standing at the bottom. Their vertical distance is 4.5 m and the angle of elevation of paul from lily is 20 degrees. Find the horizontal distance between them


1
Expert's answer
2022-06-10T11:50:03-0400

Let Paul is standing at point P, Lily is standing at point L.

Construct a right triangle "\triangle POL" ("\angle O = 90\degree"):




Vertical distance of the escalator 4.5 m, so "PO=4.5\,m".

The angle of elevation is 20 degrees, so "\angle L=20\degree".


You need to find the horizontally distance between them, so You need to find "LO".





Using the trigonometric properties of the acute angle of a right triangle we have: "\tan L=\frac{PO}{LO}".

So "LO=\frac{PO}{\tan L}=\frac{4.5\,m}{\tan20\degree}=\frac{4.5\,m}{0.3639}=12.37\,m".